∫x^4/(x^4-16)dx

Grade:12

1 Answers

Aakash
8 years ago
\int \frac{^{x^{4}}}{(x^{4}-1)}
\int \frac{^{x^{4}+1-1}}{(x^{4}-1)}
\int 1 dx -\int \frac{^{1}}{(x^{4}-1)}dx
simplify 2nd term by partial fraction method 
\int ( \frac{A}{x-1} +\frac{B}{x+1} + \frac{Cx+D}{x^{2}+1})dx
A=\frac{1}{4}
B=\frac{-1}{4}
C=0
D=\frac{-1}{2}
Now above problem reduce to standard form and Do it yourself
PLEASE APPROVE IF USEFUL